OpenStudy (anonymous):

How much work is accomplished when a force of 250 N pushes a box across the floor for a distance of 50 meters? A. 5 J B. 12,500 J C. 5 N D. 12,500 N

OpenStudy (anonymous):

the first thing you should realize here is what the units of work are. Work is equal to Force*Distance. You are given a force and you are given a distance. Dimensionally speaking it wouldn't make any since for the answer to have nothing but units of Newtons so that takes out two of your choices right there.

OpenStudy (anonymous):

Joules are not part of the answers because your force is newtons and then you multiply that by the distance that the force was applied. So D is the answer...

OpenStudy (shamim):

Required formula is W=F*d=250*50 Joule=?

OpenStudy (anonymous):

the unit of work is joules.. newton is the unit of force..so eliminate c and d.. now work=f*d =250*50=12500j option b

OpenStudy (anonymous):

b